OpenStudy (anonymous):

Show that roots of the equation x^2 + (1 - k)x = k are real for all values of k.

OpenStudy (anonymous):

You are on the right track. the discriminant = 2²-4(k-1)[-(k-3)] greater than 0 It becomes 4 + 4(k-1)(k-3) .>= 0 4(k-1)(k-3) .>= -4 (k-1)(k-3) .>= -1 k²-4k+3 >= -1 k²-4k+ 4 >= 0 (k-2)² >= 0 So k can be any real # for (k-2)² >= 0

OpenStudy (anonymous):

OK, will try and process it! Thx, Rohangrr. (:

OpenStudy (anonymous):

u call mr ron rohan or eric

OpenStudy (hoblos):

Δ = (1 - k)² -4(-k)(1) = 1-2k +k² +4k = 1+2k+k² = (1+k)² ≥ 0 for any value of k so the roots are real

OpenStudy (anonymous):

Thank youuuu Hoblos (:

OpenStudy (hoblos):

yw :)